# GAProtoNet: A Multi-head Graph Attention-based Prototypical Network for Interpretable Text Classification

### v1

@@ Abstract
Pretrained transformer-based Language Models (LMs) are well-known for their ability to achieve significant improvement on text classification tasks with their powerful word embeddings, but their black-box nature, which leads to a lack of interpretability, has been a major concern. In this work, we introduce GAProtoNet, a novel white-box Multi-head Graph Attention-based Prototypical Network designed to explain the decisions of text classification models built with LM encoders. In our approach, the input vector and prototypes are regarded as nodes within a graph, and we utilize multi-head graph attention to selectively construct edges between the input node and prototype nodes to learn an interpretable prototypical representation. During inference, the model makes decisions based on a linear combination of activated prototypes weighted by the attention score assigned for each prototype, allowing its choices to be transparently explained by the attention weights and the prototypes projected into the closest matching training examples. Experiments on multiple public datasets show our approach achieves superior results without sacrificing the accuracy of the original black-box LMs. We also compare with four alternative prototypical network variations and our approach achieves the best accuracy and F1 among all. Our case study and visualization of prototype clusters also demonstrate the efficiency in explaining the decisions of black-box models built with LMs.
@@ 1 Introduction
Deep learning models, especially transformer-based Language Models (LMs) such as BERT Devlin et al. (2018), RoBERTa (Liu et al., 2019) have significantly contributed to advancements in NLP, offering word embeddings as powerful tools for text classification. However, despite their state-of-art performance, their complexity and black-box nature obscure the decision-making process and hinder their interpretability. Now more and more real-world applications also desire classification models built with LMs to be interpretable, since it allows end-users to comprehend the decision-making process, fostering trust and encouraging adoption. To address this need, there is growing interest in enhancing the interpretability of LMs and the related text classification models in general.
@@ 2.3 Graph Attention Mechanism
For each text embedding input , we utilize matrix to transform it into query node . For each prototype , we utilized matrix to transform it into :
@@ 2.5 Training Objective
For the task of text classification, we employ the cross-entropy loss as the primary objective function to guide the model training. This loss measures the discrepancy between the predicted probability distribution from the model and the true labels, thus optimizing the model to assign higher probabilities to the correct classes. Specifically, the accuracy loss is defined as:
where is the number of samples, is the true label of sample , is the predicted probabilities.
@@
To ensure that each prototype learned by the model can find a most similar sample in the training data, we introduce the Proximity Loss. This loss measures the distance between prototypes and data points, penalizing prototype-sample pairs that are far apart. We use Euclidean distance as the distance metric and calculate the minimum distance between each prototype and all samples. The average of the minimum distances across all prototypes is then taken as the Proximity Loss:
@@ 3.1 Datasets and Metric
We evaluate our approach on three binary public benchmark datasets: Hotel Reviews 111 https://www.kaggle.com/datafiniti/ hotel-reviews.prepared by (Hong et al., 2020), IMDb 222https://huggingface.co/datasets/stanfordnlp/imdb. and Yelp Polarity Reviews. All of them are balanced datasets. To demonstrate our model’s performance on more challenging tasks, we also evaluate it on Tweet Mohammad et al. (2018) and Yahoo Zhang et al. (2015), which are 4-class and 10-class datasets seperately.
@@ 5 Related Work
Li et al. (2018) and Chen et al. (2019) introduce prototype-based architecture into interpretable image classification by proposing a prototype layer, where prototypes are randomly initialized and are made meaningful through training. Ming et al. (2019) brought this approach into NLP domain by adding a sequence encoder before the prototype layer. However, despite this intrinsically interpretable model demonstrating compelling results, there are still some performance gaps compared to the original black-box model.
@@ 6 Conclusion
We contribute an interpretable prototypical deep learning architecture that is advanced with graph attention on binary text classification. Our experiments show our approach outperforms both other interpretable and black-box architectures on three datasets. We also conduct a comprehensive analysis to show good interpretability of our approach.

### v2

@@ Abstract
Pretrained transformer-based Language Models (LMs) are well-known for their ability to achieve significant improvement on text classification tasks with their powerful word embeddings, but their black-box nature, which leads to a lack of interpretability, has been a major concern. In this work, we introduce GAProtoNet, a novel white-box Multi-head Graph Attention-based Prototypical Network designed to explain the decisions of text classification models built with LM encoders. In our approach, the input vector and prototypes are regarded as nodes within a graph, and we utilize multi-head graph attention to selectively construct edges between the input node and prototype nodes to learn an interpretable prototypical representation. During inference, the model makes decisions based on a linear combination of activated prototypes weighted by the attention score assigned for each prototype, allowing its choices to be transparently explained by the attention weights and the prototypes. Experiments on multiple public datasets show our approach achieves superior results without sacrificing the accuracy of the original black-box LMs. We also compare with four alternative prototypical network variations and our approach achieves the best accuracy and F1 among all. Our case study and visualization of prototype clusters also demonstrate the efficiency in explaining the decisions of black-box models built with LMs. Our codes are available at https://github.com/ximingwen/GAProtoNet.
@@ 1 Introduction
Deep learning models, especially transformer-based Language Models (LMs) such as BERT Devlin et al. (2018), RoBERTa (Liu et al., 2019) have significantly contributed to advancements in NLP, especially text classification. However, despite their state-of-art performance, their complexity and black-box nature obscure the decision-making process and hinder their interpretability. Now more and more real-world applications also desire classification models built with LMs to be interpretable, since it allows end-users to comprehend the decision-making process, fostering trust and encouraging adoption. To address this need, there is growing interest in enhancing the interpretability of text classification models built with LMs.
@@ 2.3 Graph Attention Mechanism
For each text embedding input , we utilize matrix to transform it into query vector . For each prototype , we utilized matrix to transform it into :
@@ 2.5 Training Objective
We employ the cross entropy loss to guide the model training. This loss measures the discrepancy between the predicted probability distribution and the true labels, optimizing the model to assign higher probabilities to the correct classes. Specifically, the accuracy loss is defined as:
where is the number of samples, while is the number of classes. is the true one-hot label of sample for class , while is the predicted probabilities.
@@
To ensure that each prototype can be projected onto a similar sample in the training data, we introduce the Proximity Loss. This loss measures the distance between prototypes and data points, penalizing prototype-sample pairs that are far apart. We use Euclidean distance as the distance metric and calculate the minimum distance between each prototype and all samples. The average of the minimum distances across all prototypes is then taken as the Proximity Loss:
@@ 3.1 Datasets and Metric
We evaluate our approach on three binary public benchmark datasets: Hotel Reviews 111 https://www.kaggle.com/datafiniti/ hotel-reviews.prepared by (Hong et al., 2020), IMDb 222https://huggingface.co/datasets/stanfordnlp/imdb. (Maas et al., 2011) and Yelp Polarity Reviews. All of them are balanced datasets. To demonstrate our model’s performance on more challenging tasks, we also evaluate it on Tweet Mohammad et al. (2018) and Yahoo Zhang et al. (2015), which are 4-class and 10-class datasets seperately.
@@ 5 Related Work
Li et al. (2018) and Chen et al. (2019) introduce prototype-based architecture into interpretable image classification by proposing a prototype layer, where prototypes are randomly initialized and are made meaningful through training. Ming et al. (2019) brought this approach into NLP domain by adding a sequence encoder before the prototype layer. Concept representation learning is also applied in enhancing the interpretability of the structure Jin et al. (2024c, b, a) while others improve interpretability by extracting minimal sentiment triplets Sun et al. (2024b, a). However, despite this intrinsically interpretable model demonstrating compelling results, there are still some performance gaps compared to the original black-box model.
@@ 6 Conclusion
We contribute an interpretable prototypical deep learning architecture that is advanced with graph attention on text classification. Our experiments show our approach outperforms both other interpretable and black-box architectures on benchmark datasets. We also conduct a comprehensive analysis to show good interpretability of our approach.